OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x - 6) and (x - 6)(x - 6)

OpenStudy (hoblos):

(x + 6)(x - 6) = x^2 - 6^2 difference of two squares (x - 6)(x - 6) = (x-6)^2 = x^2 -12x +36